# Extreme Score Distributions in Countable-Outcome Round-Robin Tournaments of Equally Strong Players

### v1

@@ 1 Introduction and Background
Another eminent mathematician, Ernst Zermelo, referred to Landau’s 1914 work (Landau, 1914) and proposed a probabilistic model in which the probability that a player wins a game is given by the ratio of that player’s strength to the sum of the strengths of the two competing players. He estimated the model parameters using the method of maximum likelihood and applied this approach to the analysis of the 1924 New York chess tournament (Zermelo, 1929). The Zermelo model was later rediscovered by Bradley and Terry (Bradley and Terry, 1952) and by Ford (Ford, 1957); see also the detailed historical discussion in (David and Edwards, 2001), as well as the probabilistic aspects of rating models discussed in (Aldous, 2017).
@@ Appendix C Proof of Proposition 2
Acknowledgement
This research is supported in part by BSF grant 2020063.

### v2

@@ 1 Introduction and Background
Another eminent mathematician, Ernst Zermelo, referred to Landau’s 1914 work (Landau, 1914) and proposed a probabilistic model in which the probability that a player wins a game is given by the ratio of that player’s strength to the sum of the strengths of the two competing players. He estimated the model parameters using the method of maximum likelihood and applied this approach to the analysis of the 1924 New York chess tournament (Zermelo, 1929). The Zermelo model was later rediscovered by Bradley and Terry (1952) and by Ford (1957); see also the detailed historical discussion in David and Edwards (2001), as well as the probabilistic aspects of rating models discussed in Aldous (2017).
@@ 3 Main Results
4 Numerics
We illustrate Theorem 1 for the chess round–robin tournament model where $X_{ij}\in\left\{0,\tfrac{1}{2},1\right\}$ with $\mathbb{P}(X_{ij}=\tfrac{1}{2})=p,\,\,\,\ \mathbb{P}(X_{ij}=1)=\mathbb{P}(X_{ji}=0)=\frac{1-p}{2},$ where $p\in\left\{1/6,\,1/3,\,2/3\right\}$. For each unordered pair $\{i,j\}$, the random vector $(X_{ij},X_{ji})$ is generated independently subject to $X_{ij}+X_{ji}=1$. For each value of $n\in\left\{10,\,100,\,500,\,1000,\,10000,\,25000,\,50000,\,100000\right\},$    $p\in\left\{1/6,\,1/3,\,2/3\right\},$ and $t\in\left\{-1,\,0\right\},$ we simulated the entire tournament, computed the normalized scores $s_{1}^{\star}(n),\ldots,s_{n}^{\star}(n)$, and formed $W_{n}(t)=\sum_{i=1}^{n}\mathbf{1}\{s_{i}^{\star}(n)>x_{n}(t)\},$ where $x_{n}(t)$ is defined in (8). We repeated the entire procedure $MC$ times, obtained an empirical probability mass function of $W_{n}(t)$, and used it to estimate the total variation distance
$\widehat{d_{\mathrm{TV}}}\!\left(\mathcal{L}(W_{n}(t)),\mathrm{Poisson}(\mathrm{e}^{-t})\right)=\frac{1}{2}\sum_{k\geq 0}\bigl|\widehat{\mathbb{P}}(W_{n}(t)=k)-\mathbb{P}(\mathrm{Poisson}(\mathrm{e}^{-t})=k)\bigr|,$
where $\widehat{\mathbb{P}}$ denotes the Monte–Carlo frequencies. The resulting values are reported in Table 1.
It is important to note that each simulation required the generation and storage of an $n\times n$ matrix of match outcomes. For $n>10000$ this required the use of high-performance computing resources with very large memory. For example, for $n=100000$ a single tournament realization required approximately $580$ GB of memory, and $500$ Monte-Carlo replications required about $22$ hours of computation.
Although the accessible values of $n$ are restricted by computational constraints, the results are consistent with the convergence rate asserted in Theorem 1.
Table 1: Empirical total variation distance between the simulated law of $W_{n}(t)$ and its Poisson limit for the chess model with p=1/6, 1/3, 2/3p=1/6,\,1/3,\,2/3. Results are shown for two regimes: t=1t=-1, and $t=0$. MCMC denotes the number of Monte–Carlo replications.
t=1t=-1 $t=0$
$n$ $\dfrac{(\log\log n)^{2}}{\log n}$ MCMC $p=\tfrac{1}{6}$ $p=\tfrac{1}{3}$ $p=\tfrac{2}{3}$ $p=\tfrac{1}{6}$ $p=\tfrac{1}{3}$ $p=\tfrac{2}{3}$
10 0.3021 10000 0.3042 0.3821 0.2915 0.1612 0.1223 0.1362
100 0.5064 10000 0.1706 0.1622 0.1206 0.0477 0.0749 0.0254
500 0.5371 10000 0.1360 0.1012 0.0943 0.0476 0.0615 0.0350
1000 0.5407 10000 0.1597 0.1415 0.0932 0.0222 0.0396 0.0159
10000 0.5353 10000 0.0955 0.1121 0.1186 0.0310 0.0350 0.0228
25000 0.5293 1000 0.0994 0.0979 0.1008 0.0403 0.0238 0.0156
50000 0.5241 1000 0.1028 0.0848 0.1278 0.0497 0.0216 0.0322
100000 0.5186 500 0.1086 0.1166 0.1092 0.0341 0.0236 0.0716
@@ Appendix C Proof of Proposition 2
Acknowledgements
We thank Boris Alemi for executing the MATLAB program on the department’s high-performance computer. This research is supported in part by BSF grant 2020063.